# CoreEditor: Consistent 3D Editing via Correspondence-constrained Diffusion

### v1

@@ IV-D Ablation Study
TABLE II: Effect of the selective editing pipeline.
@@
GaussCtrl [1] 0.257 0.128 0.372
[1] + RA (Automatic) 0.258 0.129 0.357
[1] + RA (Manual) 0.260 0.132 0.352
Ours w/o RA 0.258 0.133 0.292
Ours (Automatic) 0.267 0.145 0.283
Ours (Manual) 0.270 0.145 0.281
Figure 9: Qualitative comparison with recent methods [9, 1, 29] on prompts requiring shape changes.
@@
TABLE III: Effect of the Correspondence-constraint Attention.
@@
w/o CAA 0.250 0.124 0.378
w/o Geometric 0.245 0.118 0.351
w/o Semantic 0.266 0.128 0.294
Ours 0.270 0.145 0.281
TABLE IV: Effect of varying $\beta$.
@@
w/o $\beta$ 0.257 0.133 0.297
$\beta=0.8$ 0.265 0.136 0.292
$\beta=0.85$ 0.271 0.142 0.289
$\beta=0.9$ 0.270 0.145 0.281
$\beta=0.95$ 0.266 0.137 0.290
@@
IV-E Limitations
Similar to recent methods [1, 40, 8, 29], CoreEditor leverages the original scene geometry to improve 3D consistency. As a result, our approach is limited in its ability to significantly modify scene geometry. As illustrated by the failure case in the first row of Fig. 9, both CoreEditor and other methods fail to transform the bear statue into a car. However, unlike methods such as [40, 8], which directly warp pixels based on depth maps, CoreEditor enforces correspondence-based constraints within the diffusion latent space, where the minimal patch size is $8\times 8$. This design makes CoreEditor more robust to depth errors and allows it to handle target edits with shapes that differ from the original models. The second row of Fig. 9 demonstrates that, for prompts requiring shape modification, our method surpasses baseline methods [9, 1, 29] by enabling greater geometric changes while maintaining high-quality rendering, thanks to our consistent 2D edits.
@@ V Conclusion
We propose CoreEditor, a novel framework for text-driven 3D editing that significantly enhances the quality of edited 3D scenes by improving consistency during multi-view editing. At the heart of our method is the correspondence-constrained attention module, which enforces interactions between image patches that should remain consistent within the diffusion model. To handle complex scenes with the proposed attention module, we introduce a geometry and semantic co-supported strategy to extract comprehensive correspondences, ensuring robust multi-view editing. Additionally, we design a selective editing pipeline that empowers users to choose their preferred edits from multiple candidates, enabling a highly flexible and user-centric editing process. Extensive experiments on widely-used datasets demonstrate that CoreEditor achieves state-of-the-art editing performance. These results highlight its effectiveness in extending 2D diffusion models to consistent 3D editing, representing an important step toward more flexible and controllable 3D AIGC.

### v2

@@ IV-D Ablation Study
TABLE II: Effect of the selective editing pipeline. For CLIP-based metrics, we report both the mean and median values across all evaluation cases (shown as mean / median).
@@
GaussCtrl [61] 0.257 / 0.251 0.128 / 0.126 0.372
[61] + RA (Automatic) 0.258 / 0.245 0.129 / 0.128 0.357
[61] + RA (Manual) 0.260 / 0.253 0.132 / 0.130 0.352
Ours w/o RA 0.258 / 0.246 0.133 / 0.129 0.292
Ours (Automatic) 0.267 / 0.258 0.145 / 0.138 0.283
Ours (Manual) 0.270 / 0.259 0.145 / 0.141 0.281
@@
TABLE III: Effect of the Correspondence-constraint Attention. For CLIP-based metrics, we report both the mean and median values across all evaluation cases (shown as mean / median).
@@
w/o CAA 0.250 / 0.247 0.124 / 0.126 0.378
w/o Geometric 0.245 / 0.240 0.118 / 0.113 0.351
w/o Semantic 0.266 / 0.255 0.128 / 0.129 0.294
Ours 0.270 / 0.259 0.145 / 0.141 0.281
TABLE IV: Effect of varying $\beta$. For CLIP-based metrics, we report both the mean and median values across all evaluation cases (shown as mean / median).
@@
w/o $\beta$ 0.257 / 0.252 0.133 / 0.136 0.297
$\beta=0.8$ 0.265 / 0.257 0.136 / 0.140 0.292
$\beta=0.85$ 0.271 / 0.257 0.142 / 0.139 0.289
$\beta=0.9$ 0.270 / 0.259 0.145 / 0.141 0.281
$\beta=0.95$ 0.266 / 0.255 0.137 / 0.138 0.290
@@
TABLE V: Additional quantitative comparison with PDS [27] and Morpheus [62]. For CLIP-based metrics, we report both the mean and median values across all evaluation cases (shown as mean / median).
[…]
PDS [27] 0.247 / 0.220 0.092 / 0.088 -
Morpheus [62] 0.259 / 0.242 0.123 / 0.137 0.328
Ours 0.270 / 0.259 0.145 / 0.141 0.281
IV-E Additional Comparisons
In the main experiments, we compare CoreEditor with baselines that also apply the multi-view editing paradigm in a zero-shot manner. To more comprehensively evaluate our method, we include an SDS-based method [27] and Morpheus [62] as baselines in Tab. V and Fig. 9.
SDS-based methods often suffer from slow convergence and unstable optimization, resulting in noticeably lower rendering quality and significantly longer optimization time in practice. In contrast, Morpheus adopts a fundamentally different paradigm by finetuning a 2D diffusion model with RGB-D supervision to enhance geometry-aware editing. Although this strategy can induce more pronounced geometric changes than other baselines [61, 9], Morpheus directly edits depth through the diffusion process, which introduces strong geometric perturbations and often leads to degenerate 3D editing results. In addition, it finetunes a text-to-image diffusion model on relatively small-scale data, which disrupts the learned generative priors and leads to unstable editing behavior. Compared with these approaches, CoreEditor achieves a more favorable balance between editing strength, rendering quality, and robustness, delivering consistent high-quality results.
IV-F Generalization to Different 2D Editors
To demonstrate that the proposed CCA can generalize to different 2D editors, we integrate it into another commonly used 2D editor, InstructPix2Pix [3]. As shown in Fig. 10, incorporating CCA leads to improved multi-view consistency in the outputs.
IV-G Failure Cases and Limitations
Similar to recent methods [61, 13, 41, 29], CoreEditor leverages the original scene geometry to ensure consistent 3D editing. As a result, our approach is limited in its ability to significantly modify scene geometry. As illustrated by the failure case in the first row of Fig. 11, both CoreEditor and other methods fail to transform the bear statue into a car. However, the second row of Fig. 11 demonstrates that, for prompts requiring shape modification, our method surpasses baseline methods [9, 61, 29] by enabling greater geometric changes while maintaining high-quality rendering, thanks to our consistent 2D edits.
The stronger capability of CoreEditor to induce geometric changes compared to prior methods can be attributed to three key factors. (1) Unlike methods such as [13, 41], which directly warp pixels based on depth maps, CoreEditor enforces correspondence-based constraints within the diffusion latent space, where the minimal patch size is $8\times 8$. This design makes CoreEditor more robust to depth inaccuracies and enables it to handle target edits whose geometric shapes deviate from the original input. (2) Prompts that require geometric editing are often highly ambiguous, leading different views to produce diverse and sometimes conflicting editing results when processed independently. Existing multi-view editing methods [61] tend to implicitly average these inconsistent outcomes across views, which suppresses distinct geometric changes and results in conservative edits. In contrast, our selective editing pipeline injects the most representative editing mode through the RA module, avoiding undesired averaging effects and facilitating more pronounced and coherent geometric transformations. (3) In addition, the proposed semantic correspondence allows CCA to better associate semantically similar regions across different views, enabling more consistent feature alignment during editing. By explicitly linking corresponding semantic regions, CCA effectively propagates geometry-related changes across views, leading to more stable and coherent geometric editing results.
As further evidence, Fig. 12 compares CoreEditor with GaussCtrl [61] and our ablation variants on geometric editing, showing that both RA and semantic correspondence are crucial for inducing geometric deformations. Removing either component noticeably weakens the extent of geometric changes, while their combination enables CoreEditor to achieve stronger geometric editing results than existing methods.
@@ V Conclusion
We propose CoreEditor, a novel framework for text-driven 3D editing that significantly enhances the quality of edited 3D scenes by improving consistency during multi-view editing. At the heart of our method is the correspondence-constrained attention module, which enforces interactions between image patches that should remain consistent within the diffusion model. To handle complex scenes with the proposed attention module, we introduce a geometric and semantic co-supported strategy to extract comprehensive correspondences, ensuring robust multi-view editing. Additionally, we design a selective editing pipeline that empowers users to choose their preferred edits from multiple candidates, enabling a highly flexible and user-centric editing process. Extensive experiments on widely-used datasets demonstrate that CoreEditor achieves state-of-the-art editing performance, offering superior quality and adaptability compared to existing methods. While the current implementation of CCA is designed for U-Net–based 2D diffusion editors, extending it to architectures that incorporate explicit positional encoding within the attention mechanism (e.g., FLUX [28]) would require non-trivial architectural adaptations to jointly account for positional information and cross-view correspondence. We leave this direction for future work.
VI Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. T2322012, No. 62572240, 62172218), and the Shenzhen Science and Technology Program (No. JCYJ20220818103401003, No. JCYJ20220530172403007).